1a You have a fundamental standing wave at a frequency f, tension F, linear density μ and length L. What tension would you use to double the frequency?

                                         
 

1b. You have a fundamental standing wave at a frequency f, tension F, linear density μ and length L. What length would you use to double the frequency?

 

1c. You have a fundamental standing wave at a frequency f, tension F, linear density μ and length L. What linear density would you use to double the frequency?

 

1d. You have a fundamental standing wave at a frequency of 500 Hz, tension F, linear density μ and length L. What would the new fundamental frequency be if you doubled the tension, and at the same time cut the length in half and cut the linear density in half?

ALSO I WOULD REALLY APPRECIATE AN EXPLANATION, THANKS.
Answer #1

Solution 1)a)

We know,

f = 1 / 2L √ F / μ

So, from the above equation, we have

f1 / f2 = √ F1 / F2

or ( f1 / f2)2= F1 / F2

(f1 / 2 f1 )2 = F1 / F2

1/ 4 = F1 /F2

or F2 = 4 F1 (Ans)

=======

Acc to HOMEWORKLIB RULES, I can answer only one at a time. Good luck!:)